# The effect of priors on Learning with Restricted Boltzmann Machines

### v1

@@ Appendix A Derivation of the RS equations
$\sum_{\bm{\mathcal{S}}}$ $\sum_{\bm{\xi}^{1},\ldots,\bm{\xi}^{n}}\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\mathbb{E}_{\hat{\bm{\xi}}}\;\mathbb{E}_{\hat{\bm{\tau}}}e^{\sum_{i,a}{s}_{i}^{a}\left(\sqrt{\frac{\hat{\beta}}{N}}\hat{\tau}_{a}\hat{\xi}_{i}+\sqrt{\frac{\beta}{N}}\sum_{\gamma}\tau_{a}^{\gamma}\xi_{i}^{\gamma}\right)}=\sum_{\bm{\xi}^{1},\ldots,\bm{\xi}^{n}}\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\mathbb{E}_{\hat{\bm{\xi}}}\;\mathbb{E}_{\hat{\bm{\tau}}}e^{\sum_{i,a}\ln\cosh(\left(\sqrt{\frac{\hat{\beta}}{N}}\hat{\tau}_{a}\hat{\xi}_{i}+\sqrt{\frac{\beta}{N}}\sum_{\gamma}\tau_{a}^{\gamma}\xi_{i}^{\gamma}\right)}$
$\approx$ $\sum_{\bm{\xi}^{1},\ldots,\bm{\xi}^{n}}\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\;\mathbb{E}_{\hat{\bm{\xi}}}\;\mathbb{E}_{\hat{\bm{\tau}}}\,e^{\frac{M\hat{\beta}}{2}d_{\hat{\tau}}d_{\hat{\xi}}+M\beta\sum_{\gamma<\gamma^{\prime}}q_{\tau}^{\gamma\gamma^{\prime}}q^{\gamma\gamma^{\prime}}+\frac{M\beta}{2}\sum_{\gamma}d^{\gamma}_{\tau}d^{\gamma}+M\sqrt{\hat{\beta}\beta}\sum_{\gamma}m_{\tau}^{\gamma}m^{\gamma}}\;,$ (84)
@@
Consequently, we can fix the values of these parameters, as previously done in (83), and incorporate the corresponding densities of states $\mathcal{D}(\bm{\Lambda},\bm{\Lambda}_{\tau}):=\mathcal{D}(\bm{m},\bm{q},\bm{d})\,\mathcal{D}(\bm{m}_{\tau},\bm{q}_{\tau},\bm{d}_{\tau})\,\mathcal{D}(d_{\hat{\xi}},d_{\hat{\tau}})$. Both the densities affect the extremized functions $f_{T}$ and $f_{S}$, as their parameters depend respectively on $d_{\hat{\xi}}$ and $d^{\gamma}$. For brevity of notation we defined the sets of order parameters: $\bm{\Lambda}:=\{\bm{p},\bm{m},\bm{q},\bm{d},d_{\hat{\xi}}\}$ and $\bm{\Lambda}_{\tau}:=\{\bm{p}_{\tau},\bm{m}_{\tau},\bm{q}_{\tau},\bm{d}_{\tau},d_{\hat{\tau}}\}$ and their conjugated ones $\hat{\bm{\Lambda}}$, $\bm{\hat{\Lambda}}_{\tau}$, obtaining the following form of the averaged partition function
@@
$+$ $\frac{M\beta}{2}\sum_{\gamma}d^{\gamma}_{\tau}d^{\gamma}+M\sqrt{\hat{\beta}\beta}\sum_{\gamma}m_{\tau}^{\gamma}m^{\gamma}+N\hat{\beta}\operatorname{\mathrm{Extr}}_{\bm{p}_{T}}f_{T}(\bm{p}_{T})+N\beta\operatorname{\mathrm{Extr}}_{\bm{p}_{S}}f_{S}(\bm{p}_{S})\Big{\}}\,.$
@@
$-\beta f\approx$ $\lim_{\begin{subarray}{c}N\to\infty\\ n\to 0\end{subarray}}\frac{1}{Nn}\ln\Bigg{(}e^{N\operatorname{\mathrm{Extr}}_{\begin{subarray}{c}\scriptstyle\bm{\Lambda},\bm{\Lambda}_{\tau}\\ \scriptstyle\hat{\bm{\Lambda}},\hat{\bm{\Lambda}}_{\tau}\end{subarray}}\hat{f}(\bm{\Lambda},\bm{\Lambda}_{\tau},\hat{\bm{\Lambda}},\hat{\bm{\Lambda}}_{\tau})}\Bigg{)}\approx\lim_{n\to 0}\frac{1}{n}\operatorname{\mathrm{Extr}}_{\begin{subarray}{c}\scriptstyle\bm{\Lambda},\bm{\Lambda}_{\tau}\\ \scriptstyle\hat{\bm{\Lambda}},\hat{\bm{\Lambda}}_{\tau}\end{subarray}}\hat{f}(\bm{\Lambda},\bm{\Lambda}_{\tau},\hat{\bm{\Lambda}},\hat{\bm{\Lambda}}_{\tau})$ (86)
$-\beta\hat{f}(\bm{\Lambda},\bm{\Lambda}_{\tau},\hat{\bm{\Lambda}},\hat{\bm{\Lambda}}_{\tau})$ $=\frac{\alpha\hat{\beta}}{2}d_{\hat{\tau}}d_{\hat{\xi}}+\alpha\beta\sum_{\gamma<\gamma^{\prime}}q_{\tau}^{\gamma\gamma^{\prime}}q^{\gamma\gamma^{\prime}}+\frac{\beta}{2}\alpha^{2}\Omega_{\xi}\sum_{\gamma}(p_{\tau}^{\gamma})^{2}+\frac{\beta}{2}\Omega_{\tau}\sum_{\gamma}(p^{\gamma})^{2}+$ (87)
@@
$+\alpha\log\mathbb{E}_{\hat{\tau}}\mathbb{E}_{\tilde{s}}\mathbb{E}_{\{\tau^{\gamma}\}}\exp{\left(\hat{d}_{\hat{\tau}}\,\hat{\tau}^{2}+\frac{1}{2}\sum_{\gamma\gamma^{\prime}}\hat{q}_{\tau}^{\gamma\gamma^{\prime}}\tau^{\gamma}\tau^{\gamma^{\prime}}-\frac{1}{2}\sum_{\gamma}\hat{q}_{\tau}^{\gamma\gamma}(\tau^{\gamma})^{2}+\sum_{\gamma}\hat{m}_{\tau}^{\gamma}\tau^{\gamma}\hat{\tau}+\sum_{\gamma}\hat{d}_{\tau}^{\gamma}(\tau^{\gamma})^{2}+\sum_{\gamma}\tilde{s}p_{\tau}^{\gamma}\tau^{\gamma}\right)}+$
$+\log\mathbb{E}_{\hat{\xi}}\mathbb{E}_{s}\mathbb{E}_{\{{\xi}^{\gamma}\}}\exp{\left(\hat{d}_{\hat{\xi}}\,\hat{\xi}^{2}+\frac{1}{2}\sum_{\gamma\gamma^{\prime}}\hat{q}^{\gamma\gamma^{\prime}}\xi^{\gamma}\xi^{\gamma^{\prime}}-\frac{1}{2}\sum_{\gamma}\hat{q}^{\gamma\gamma}(\xi^{\gamma})^{2}+\sum_{\gamma}\hat{m}^{\gamma}\xi^{\gamma}\hat{\xi}+\sum_{\gamma}\hat{d}^{\gamma}(\xi^{\gamma})^{2}+\sum_{\gamma}s\,\hat{p}^{\gamma}\xi^{\gamma}\right)}\,,$
@@
$=\log\left[\mathbb{E}_{\hat{\xi}}\mathbb{E}_{s}\mathbb{E}_{\{{\xi}^{\gamma}\}}\exp{\left(\hat{d}_{\hat{\xi}}\hat{\xi}^{2}+\frac{1}{2}\sum_{\gamma\gamma^{\prime}}\hat{q}^{\gamma\gamma^{\prime}}\xi^{\gamma}\xi^{\gamma^{\prime}}-\frac{1}{2}\sum_{\gamma}\hat{q}^{\gamma\gamma}(\xi^{\gamma})^{2}+\sum_{\gamma}\hat{m}^{\gamma}\xi^{\gamma}\hat{\xi}+\sum_{\gamma}\hat{d}^{\gamma}(\xi^{\gamma})^{2}+\sum_{\gamma}s\,\hat{p}^{\gamma}\xi^{\gamma}\right)}\right]^{N}\,,$
@@
$\mathbb{E}_{\hat{\tau}}e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}\mathbb{E}_{\tilde{s}}\,\mathbb{E}_{\{\tau^{\gamma}\}}\exp{\left(\frac{1}{2}\hat{q}_{\tau}\left(\sum_{\gamma}\tau^{\gamma}\right)^{2}-\frac{1}{2}\hat{q}_{\tau}\sum_{\gamma}(\tau^{\gamma})^{2}+\hat{m}_{\tau}\sum_{\gamma}\tau^{\gamma}\hat{\tau}+\hat{d}_{\tau}\sum_{\gamma}(\tau^{\gamma})^{2}+\tilde{s}\hat{p}_{\tau}\sum_{\gamma}\tau^{\gamma}\right)}=$
$=$ $\Bigg{\langle}\mathbb{E}_{\hat{\tau}}\,e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}\mathbb{E}_{\tilde{s}}\Bigg{(}\mathbb{E}_{\tau}\exp\left(-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\tilde{s}\,\hat{p}_{\tau})\right)\Bigg{)}^{n}\Bigg{\rangle}_{z}$
$=$ $\Bigg{\langle}\mathbb{E}_{\hat{\tau}}\,e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}\mathbb{E}_{\tilde{s}}\exp\left\{n\log\Bigg{[}\mathbb{E}_{\tau}\exp\left(-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\tilde{s}\,\hat{p}_{\tau})\right)\Bigg{]}\right\}\Bigg{\rangle}_{z}$
$\simeq$ $\Bigg{\langle}\mathbb{E}_{\hat{\tau}}\,e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}\mathbb{E}_{\tilde{s}}\left\{1+n\log\Bigg{[}\mathbb{E}_{\tau}\exp\left(-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\tilde{s}\,\hat{p}_{\tau})\right)\Bigg{]}\right\}\Bigg{\rangle}_{z}$
$=$ $\Bigg{\langle}\mathbb{E}_{\hat{\tau}}\,e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}+n\mathbb{E}_{\hat{\tau}}\,e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}\mathbb{E}_{\tilde{s}}\log\Bigg{[}\mathbb{E}_{\tau}\exp\left(-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\tilde{s}\,\hat{p}_{\tau})\right)\Bigg{]}\Bigg{\rangle}_{z}$
$=$ $\mathbb{E}_{\hat{\tau}}\,e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}\,\left\{1+\frac{n}{\mathbb{E}_{\hat{\tau}}\,e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}}\Bigg{\langle}\mathbb{E}_{\hat{\tau}}\,e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}\mathbb{E}_{\tilde{s}}\log\Bigg{[}\mathbb{E}_{\tau}\exp\left(-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\tilde{s}\,\hat{p}_{\tau})\right)\Bigg{]}\Bigg{\rangle}_{z}\right\}\,,$
@@
$\alpha\log\left(\mathbb{E}_{\hat{\tau}}\,e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}\,\left\{1+\frac{n}{\mathbb{E}_{\hat{\tau}}\,e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}}\Bigg{\langle}\mathbb{E}_{\hat{\tau}}\,e^{\hat{d}_{\hat{\tau}}\hat{\tau}^{2}}\mathbb{E}_{\tilde{s}}\log\mathbb{E}_{\tau}\exp\left(-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\tilde{s}\,\hat{p}_{\tau})\right)\Bigg{\rangle}_{z}\right\}\right)\simeq$
@@
$\log\mathbb{E}_{\hat{{\xi}}}e^{\hat{d}_{\hat{\xi}}\hat{\xi}^{2}}\mathbb{E}_{s}\,\mathbb{E}_{\{{\xi}^{\gamma}\}}\exp{\left(\frac{\hat{q}}{2}\bigg{(}\sum_{\gamma}\xi^{\gamma}\bigg{)}^{2}-\frac{\hat{q}}{2}\sum_{\gamma}(\xi^{\gamma})^{2}+\hat{m}\sum_{\gamma}\xi^{\gamma}\hat{\xi}+\hat{d}\sum_{\gamma}(\xi^{\gamma})^{2}+s\,\hat{p}\sum_{\gamma}\xi^{\gamma}\right)}\simeq$
@@
$-\beta f^{RS}(\beta,\hat{\beta},\gamma)\approx\lim_{n\to 0}\frac{1}{n}\operatorname{\mathrm{Extr}}_{\tiny\begin{subarray}{c}\Lambda,\hat{\Lambda}\\ \tiny\Lambda_{\tau},\hat{\Lambda}_{\tau}\end{subarray}}f^{RS}(\Lambda,\hat{\Lambda},\Lambda_{\tau},\hat{\Lambda}_{\tau})\,,$

### v2

@@ Appendix A Derivation of the RS equations
$\sum_{\bm{\mathcal{S}}}\sum_{\bm{\xi}^{1},\ldots,\bm{\xi}^{n}}\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\mathbb{E}_{\hat{\bm{\xi}}}\;\mathbb{E}_{\hat{\bm{\tau}}}\exp\left(\sum_{i,a}{s}_{i}^{a}\left(\sqrt{\frac{\hat{\beta}}{N}}\hat{\tau}_{a}\hat{\xi}_{i}+\sqrt{\frac{\beta}{N}}\sum_{\gamma}\tau_{a}^{\gamma}\xi_{i}^{\gamma}\right)\right)=$ (84)
$=\sum_{\bm{\xi}^{1},\ldots,\bm{\xi}^{n}}\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\mathbb{E}_{\hat{\bm{\xi}}}\;\mathbb{E}_{\hat{\bm{\tau}}}\exp\left(\sum_{i,a}\ln\cosh(\left(\sqrt{\frac{\hat{\beta}}{N}}\hat{\tau}_{a}\hat{\xi}_{i}+\sqrt{\frac{\beta}{N}}\sum_{\gamma}\tau_{a}^{\gamma}\xi_{i}^{\gamma}\right)\right)$
$\approx\sum_{\bm{\xi}^{1},\ldots,\bm{\xi}^{n}}\mathbb{E}_{\{\bm{\tau}^{\gamma}\}}\;\mathbb{E}_{\hat{\bm{\xi}}}\;\mathbb{E}_{\hat{\bm{\tau}}}\,\exp\left(\frac{M\hat{\beta}}{2}d_{\hat{\tau}}d_{\hat{\xi}}+M\beta\sum_{\gamma<\gamma^{\prime}}q_{\tau}^{\gamma\gamma^{\prime}}q^{\gamma\gamma^{\prime}}+\frac{M\beta}{2}\sum_{\gamma}d^{\gamma}_{\tau}d^{\gamma}+M\sqrt{\hat{\beta}\beta}\sum_{\gamma}m_{\tau}^{\gamma}m^{\gamma}\right)\;,$ (85)
@@
Consequently, we can fix the values of these parameters, as previously done in (83), and incorporate the corresponding densities of states $\mathcal{D}(\bm{\Lambda},\bm{\Lambda}_{\tau}):=\mathcal{D}(\bm{m},\bm{q},\bm{d})\,\mathcal{D}(\bm{m}_{\tau},\bm{q}_{\tau},\bm{d}_{\tau})\,$
@@
$\hskip-34.14322pt\left[Z^{n}\right]^{\bm{\mathcal{S}}}=\int d\,\bm{\Lambda}d\,\bm{\Lambda}_{\tau}\mathcal{D}(\bm{\Lambda},\bm{\Lambda}_{\tau})\exp$ $\Big(\frac{M\hat{\beta}}{2}d_{\hat{\tau}}d_{\xi}+M\beta\sum_{\gamma<\gamma^{\prime}}q_{\tau}^{\gamma\gamma^{\prime}}q^{\gamma\gamma^{\prime}}+\frac{N\beta}{2}\alpha^{2}\Omega\sum_{\gamma}(p_{\tau}^{\gamma})^{2}+\frac{N\beta}{2}\Omega_{\tau}\sum_{\gamma}(p^{\gamma})^{2}+$ (86)
$+$ $\frac{M\beta}{2}\sum_{\gamma}d^{\gamma}_{\tau}d^{\gamma}+M\sqrt{\hat{\beta}\beta}\sum_{\gamma}m_{\tau}^{\gamma}m^{\gamma}+N\hat{\beta}\operatorname{\mathrm{Extr}}_{\bm{p}_{T}}f_{T}(\bm{p}_{T})+N\beta\operatorname{\mathrm{Extr}}_{\bm{p}_{S}}f_{S}(\bm{p}_{S})\Big)\,.$
@@
$-\beta f\approx$ $\lim_{\begin{subarray}{c}N\to\infty\\ n\to 0\end{subarray}}\frac{1}{Nn}\ln\Bigg(\exp\left(N\operatorname{\mathrm{Extr}}_{\begin{subarray}{c}\scriptstyle\bm{\Lambda},\bm{\Lambda}_{\tau}\\ \scriptstyle\hat{\bm{\Lambda}},\hat{\bm{\Lambda}}_{\tau}\end{subarray}}\hat{f}(\bm{\Lambda},\bm{\Lambda}_{\tau},\hat{\bm{\Lambda}},\hat{\bm{\Lambda}}_{\tau})\right)\Bigg)\approx\lim_{n\to 0}\frac{1}{n}\operatorname{\mathrm{Extr}}_{\begin{subarray}{c}\scriptstyle\bm{\Lambda},\bm{\Lambda}_{\tau}\\ \scriptstyle\hat{\bm{\Lambda}},\hat{\bm{\Lambda}}_{\tau}\end{subarray}}\hat{f}(\bm{\Lambda},\bm{\Lambda}_{\tau},\hat{\bm{\Lambda}},\hat{\bm{\Lambda}}_{\tau})$ (87)
$\hskip-14.22636pt-\beta\hat{f}(\bm{\Lambda},\bm{\Lambda}_{\tau},\hat{\bm{\Lambda}},\hat{\bm{\Lambda}}_{\tau})$ $=\frac{\alpha\hat{\beta}}{2}d_{\hat{\tau}}d_{\hat{\xi}}+\alpha\beta\sum_{\gamma<\gamma^{\prime}}q_{\tau}^{\gamma\gamma^{\prime}}q^{\gamma\gamma^{\prime}}+\frac{\beta}{2}\alpha^{2}\Omega_{\xi}\sum_{\gamma}(p_{\tau}^{\gamma})^{2}+\frac{\beta}{2}\Omega_{\tau}\sum_{\gamma}(p^{\gamma})^{2}+$ (88)
@@
$\hskip-85.35826pt+\alpha\log\mathbb{E}_{\hat{\tau}}\mathbb{E}_{\tilde{s}}\mathbb{E}_{\{\tau^{\gamma}\}}\exp{\left(\hat{d}_{\hat{\tau}}\,\hat{\tau}^{2}+\frac{1}{2}\sum_{\gamma\gamma^{\prime}}\hat{q}_{\tau}^{\gamma\gamma^{\prime}}\tau^{\gamma}\tau^{\gamma^{\prime}}-\frac{1}{2}\sum_{\gamma}\hat{q}_{\tau}^{\gamma\gamma}(\tau^{\gamma})^{2}+\sum_{\gamma}\hat{m}_{\tau}^{\gamma}\tau^{\gamma}\hat{\tau}+\sum_{\gamma}\hat{d}_{\tau}^{\gamma}(\tau^{\gamma})^{2}+\sum_{\gamma}\tilde{s}p_{\tau}^{\gamma}\tau^{\gamma}\right)}+$
$\hskip-85.35826pt+\log\mathbb{E}_{\hat{\xi}}\mathbb{E}_{s}\mathbb{E}_{\{{\xi}^{\gamma}\}}\exp{\left(\hat{d}_{\hat{\xi}}\,\hat{\xi}^{2}+\frac{1}{2}\sum_{\gamma\gamma^{\prime}}\hat{q}^{\gamma\gamma^{\prime}}\xi^{\gamma}\xi^{\gamma^{\prime}}-\frac{1}{2}\sum_{\gamma}\hat{q}^{\gamma\gamma}(\xi^{\gamma})^{2}+\sum_{\gamma}\hat{m}^{\gamma}\xi^{\gamma}\hat{\xi}+\sum_{\gamma}\hat{d}^{\gamma}(\xi^{\gamma})^{2}+\sum_{\gamma}s\,\hat{p}^{\gamma}\xi^{\gamma}\right)}\,$,
@@
$\hskip-28.45274pt=\log\left[\mathbb{E}_{\hat{\xi}}\mathbb{E}_{s}\mathbb{E}_{\{{\xi}^{\gamma}\}}\exp{\left(\hat{d}_{\hat{\xi}}\hat{\xi}^{2}+\frac{1}{2}\sum_{\gamma\gamma^{\prime}}\hat{q}^{\gamma\gamma^{\prime}}\xi^{\gamma}\xi^{\gamma^{\prime}}-\frac{1}{2}\sum_{\gamma}\hat{q}^{\gamma\gamma}(\xi^{\gamma})^{2}+\sum_{\gamma}\hat{m}^{\gamma}\xi^{\gamma}\hat{\xi}+\sum_{\gamma}\hat{d}^{\gamma}(\xi^{\gamma})^{2}+\sum_{\gamma}s\,\hat{p}^{\gamma}\xi^{\gamma}\right)}\right]^{N}\,$,
@@
$\hskip-28.45274pt\mathbb{E}_{\hat{\tau}}\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)\mathbb{E}_{\tilde{s}}\,\mathbb{E}_{\{\tau^{\gamma}\}}\exp{\left(\frac{1}{2}\hat{q}_{\tau}\left(\sum_{\gamma}\tau^{\gamma}\right)^{2}-\frac{1}{2}\hat{q}_{\tau}\sum_{\gamma}(\tau^{\gamma})^{2}+\hat{m}_{\tau}\sum_{\gamma}\tau^{\gamma}\hat{\tau}+\hat{d}_{\tau}\sum_{\gamma}(\tau^{\gamma})^{2}+\tilde{s}\hat{p}_{\tau}\sum_{\gamma}\tau^{\gamma}\right)}=$
$=$ $\Bigg\langle\mathbb{E}_{\hat{\tau}}\,\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)\mathbb{E}_{\tilde{s}}\Bigg(\mathbb{E}_{\tau}\exp\left(-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\tilde{s}\,\hat{p}_{\tau})\right)\Bigg)^{n}\Bigg\rangle_{z}$
$=$ $\Bigg\langle\mathbb{E}_{\hat{\tau}}\,\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)\mathbb{E}_{\tilde{s}}\exp\left\{n\log\Bigg[\mathbb{E}_{\tau}\exp\left(-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\tilde{s}\,\hat{p}_{\tau})\right)\Bigg]\right\}\Bigg\rangle_{z}$
$\simeq$ $\Bigg\langle\mathbb{E}_{\hat{\tau}}\,\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)\mathbb{E}_{\tilde{s}}\left\{1+n\log\Bigg[\mathbb{E}_{\tau}\exp\left(-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\tilde{s}\,\hat{p}_{\tau})\right)\Bigg]\right\}\Bigg\rangle_{z}$
$\simeq$ $\Bigg\langle\mathbb{E}_{\hat{\tau}}\,\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)+n\mathbb{E}_{\hat{\tau}}\,\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)\mathbb{E}_{\tilde{s}}\log\Bigg[\mathbb{E}_{\tau}\exp\left(-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\tilde{s}\,\hat{p}_{\tau})\right)\Bigg]\Bigg\rangle_{z}$
$\simeq\mathbb{E}_{\hat{\tau}}\,\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)\,\left\{1+\frac{n}{\mathbb{E}_{\hat{\tau}}\,\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)}\Bigg\langle\mathbb{E}_{\hat{\tau}}\,\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)\mathbb{E}_{\tilde{s}}\log\Bigg[\mathbb{E}_{\tau}\exp\left(-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\tilde{s}\,\hat{p}_{\tau})\right)\Bigg]\Bigg\rangle_{z}\right\}\,$,
@@
$\alpha\log\left(\mathbb{E}_{\hat{\tau}}\,\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)\,\left\{1+\frac{n}{\mathbb{E}_{\hat{\tau}}\,\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)}\Bigg\langle\mathbb{E}_{\hat{\tau}}\,\exp\left(\hat{d}_{\hat{\tau}}\hat{\tau}^{2}\right)\mathbb{E}_{\tilde{s}}\log\mathbb{E}_{\tau}\exp\left(-(\frac{\hat{q}_{\tau}}{2}-\hat{d}_{\tau})\tau^{2}+\tau(\hat{m}_{\tau}\hat{\tau}+\sqrt{\hat{q}_{\tau}}z+\tilde{s}\,\hat{p}_{\tau})\right)\Bigg\rangle_{z}\right\}\right)\simeq$
@@
$\log\mathbb{E}_{\hat{{\xi}}}\exp\left(\hat{d}_{\hat{\xi}}\hat{\xi}^{2}\right)\mathbb{E}_{s}\,\mathbb{E}_{\{{\xi}^{\gamma}\}}\exp{\left(\frac{\hat{q}}{2}\bigg(\sum_{\gamma}\xi^{\gamma}\bigg)^{2}-\frac{\hat{q}}{2}\sum_{\gamma}(\xi^{\gamma})^{2}+\hat{m}\sum_{\gamma}\xi^{\gamma}\hat{\xi}+\hat{d}\sum_{\gamma}(\xi^{\gamma})^{2}+s\,\hat{p}\sum_{\gamma}\xi^{\gamma}\right)}\simeq$
@@
$\beta f(\beta,\hat{\beta},\gamma)\approx\lim_{n\to 0}\frac{1}{n}\operatorname{\mathrm{Extr}}_{\tiny\begin{subarray}{c}\Lambda,\hat{\Lambda}\\ \tiny\Lambda_{\tau},\hat{\Lambda}_{\tau}\end{subarray}}\hat{f}^{RS}(\Lambda,\hat{\Lambda},\Lambda_{\tau},\hat{\Lambda}_{\tau})\,$,